If two dice are rolled one time, find the probability of getting these results.
a. A sum of 9
b. sum of 7 or 11
c. Doubles
d. A sum less than 9
e. A sum greater than or equal to 10
1
Expert's answer
2020-12-11T13:10:14-0500

When two dice are thrown simultaneously, thus number of event can be 62 = 36 because each die has 1 to 6 number on its faces.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & 1 & 2 & 3 & 4&5 & 6 \\ \hline
 1 & (1,1) & (1,2) & (1,3)& (1,4) & (1,5) & (1,6) \\
 \hdashline
 2 & (2,1) & (2,2) & (2,3)& (2,4) & (2,5) & (2,6) \\
 \hdashline
 3 & (3,1) & (3,2) & (3,3)& (3,4) & (3,5) & (3,6) \\
 \hdashline
 4 & (4,1) & (4,2) & (4,3)& (4,4) & (4,5) & (4,6) \\
 \hdashline
 5 & (5,1) & (5,2) & (5,3)& (5,4) & (5,5) & (5,6) \\
 \hdashline
6 & (6,1) & (6,2) & (6,3)& (6,4) & (6,5) & (6,6) \\
 \hdashline
\end{array}"

a. A sum of 9


"(3,6), (4,5), (5,4), (6,3)"

"P(sum=9)=\dfrac{4}{36}=\dfrac{1}{9}"

b. sum of 7 or 11


"(1,6),(2,5),(3,4), (4,3), (5,2),(5,6), (6,1),(6,5)"

"P(sum=9\ or\ sum=11)=\dfrac{8}{36}=\dfrac{2}{9}"

c. Doubles


"(1,1), (2,2), (3,3), (4,4),(5,5),(6,6)"

"P(doubles)=\dfrac{6}{36}=\dfrac{1}{6}"

d. A sum less than 9


"(1,1), (1,2), (1,3), (1,4),(1,5),(1,6),""(2,1), (2,2), (2,3), (2,4),(2,5),(2,6),""(3,1), (3,2), (3,3), (3,4),(3,5),""(4,1), (4,2), (4,3), (4,4),""(5,1), (5,2), (5,3),""(6,1), (6,2)"

"P(Sum<9)=\dfrac{26}{36}=\dfrac{13}{18}"



e. A sum greater than or equal to 10


"(4,6), (5,5), (5,6), (6,4),(6,5),(6,6)"

"P(Sum\geq10)=\dfrac{6}{36}=\dfrac{1}{6}"